# Mamba-ND: Selective State Space Modeling for Multi-Dimensional Data

### 08 Feb 2024 v1

@@ Abstract
In recent years, Transformers have become the de-facto architecture for sequence modeling on text and a variety of multi-dimensional data, such as images and video. However, the use of self-attention layers in a Transformer incurs prohibitive compute and memory complexity that scales quadratically w.r.t. the sequence length. A recent architecture, Mamba, based on state space models has been shown to achieve comparable performance for modeling text sequences, while scaling linearly with the sequence length. In this work, we present Mamba-ND, a generalized design extending the Mamba architecture to arbitrary multi-dimensional data. Our design alternatively unravels the input data across different dimensions following row-major orderings. We provide a systematic comparison of Mamba-ND with several other alternatives, based on prior multi-dimensional extensions such as Bi-directional LSTMs and S4ND. Empirically, we show that Mamba-ND demonstrates performance competitive with the state-of-the-art on a variety of multi-dimensional benchmarks, including ImageNet-1K classification, HMDB-51 action recognition, and ERA5 weather forecasting.
@@ 1 Introduction
The design of flexible and scalable neural network architectures is fundamental to the success of deep learning across diverse domains. Convolutional neural networks [23] excel at handling continuous data such as images, audio, and video. Recently, they have been surpassed by Transformers [45], which process continuous data as a discrete sequence of patches [12]. Despite their superior performance on many tasks such as Image Classification and Video Action Recognition, Transformer-based models struggle to scale to larger patch sequence as they scale quadratically with respect to sequence length. Most recently, a special kind of State Space Model (SSM) termed as Mamba [15], has demonstrated stronger performance than Transformers while maintaining a linear complexity. However, the impressive performance of Mamba was shown on 1D text sequences. This leaves open the question whether Mamba can be effectively extended to multi-dimensional data such as images, video, or scientific datasets, which is the focus of this work.
@@
Unlike convolution or self-attention operations, which can be computed in parallel across the ND input data, Mamba requires a specific ordering of the data. Determining such an order is not an easy task. Building on past work in architecture design, we could consider many choices. One naive approach is to flatten the data in row-major order. Intuitively, this is non-optimal because in this setting information only flows in one direction, which is suboptimal for modeling complex relationships over multi-dimensional data that do not have any default ordering. Drawing inspiration from early works on Bi-directional LSTM, another alternative is to process the sequence in two directions at each layer and aggregate the results. We call this method Bi-SSM. While this design, in principle, allows information exchange between two arbitrary patches, two patches adjacent to each other spatially may have a huge distance between them on the computation graph. A natural extension of this method, which we call ND-SSM, is to process the input in $2D$ directions, where $D$ is the dimension of the data, and aggregate the results.
Rather than picking orders in sequence, another possibility is to borrow inspiration from multi-head self-attention in transformers, wherein we can split the channels into multiple heads and let each head process an SSM in a different direction. This design is similar to ND-SSM but differs in that it has less computational burden within each layer.
@@
At a high level, there is also the question of block-level design, which specifies how Mamba layers are organized. For example, even using the vanilla Mamba layer as a black box, one can still apply the bi-directional or N-directional design at the block level by processing the same input through multiple layers with different directions.
@@
In this work, we conducted an extensive study on these possible design choices. Surprisingly, we find that simply alternating between three fixed row-major orderings is one of the best-performing strategies on both 2D and 3D data. Armed with these findings, we propose Mamba-ND, a surprisingly simple yet effective design to extend Mamba to multi-dimensional data. Mamba-ND does not introduce any complicated changes to 1D-SSM layers. By stacking 1D-SSM layers as black boxes and alternating the sequence order between each layer, Mamba-ND was able to surpass Transformer-based models on various tasks, including image classification, action recognition, and weather forecasting, with a lower parameter count. It also maintains a linear complexity with respect to input sequence length. We show these results in Figure 1.
@@
  • We proposed Mamba-ND , which extends SSM to higher dimension through simply alternating sequence odering across layers.
  • Compared with Transformers, our model was able to achieve stronger performance at lower parameter count. Compared with ViT, our model achieves +1.5 increase in ImageNet-1K classification accuracy while reducing the number of parameters by 20.7%. On HMDB-51 video classification task, Mamba-ND  our model achieves +0.9 increase compared with 3D Swin-Transformer while reducing the number of parameters by 39.0%. On ERA5 5.625 degree 72 hour weather forecast tasks, Mamba-ND  achieves +0.7 gain in Anomaly Correlation Coefficient (ACC) while reducing the number of parameters by 44.5% when compared with Cli-ViT.
  • We conducted extensive ablation studies on various more complicated approaches that extend SSM to higher dimension. We find that complicated designs do not necessarily translate into stronger performance. These results should help future work on adapting 1D Sequence models to higher dimension.
@@ 2 Related Works
Image Modeling.
Convolutional Neural Networks (CNNs) [23, 17, 20, 39, 21, 38] have historically been the state-of-the-art approach for image recognition tasks. These networks learn to extract hierarchical features from images by using interleaved convolution and pooling layers. While they have demonstrated strong performance, the receptive field of convolution kernels only grows linearly with the model depth. This limitation restricts the ability of CNNs to effectively capture global context.
More recently, Vision Transformers (ViTs) [12] have emerged as a strong candidate for vision tasks. Drawing inspiration from the field of natural language processing, these models first divide an input image into a grid of discrete patches and then process them as a 1D sequence. While they provide a global receptive field through self-attention mechanisms at every layer, such advantages come with a cost. In particular, they exhibit quadratic computational and memory complexity with respect to input length, making them challenging to scale to large images.
To address this limitation, several works have proposed hybrid architectures that incorporate attention mechanisms into CNNs [47, 5, 24] or introduce hierarchical designs into transformers [26, 25, 8, 10, 11, 29, 40, 46, 52]. Other approaches, such as PixelRNN [43], attempt to apply recurrent models, which are inherently designed for linear complexity and global context, to images.
There has also been considerable efforts to apply state space models to 2D images, such as 2D-SSM[2] and S4ND [33]. Our work falls within the this category, wherein we specifically explore ways to adapt selective state space models to images.
Video Modeling.
Modeling videos has long presented a significant challenge. Earlier works [4, 41] primarily focused on extending 2D ConvNets into the 3D domain. More recently, 3D Transformers [1, 28, 32] have demonstrated superior performance. Some of these works add temporal attention on top of existing 2D Vision Transformers, while others attempt to apply self-attention in a pure 3D fashion.
The unique characteristics of the temporal dimension in video data have inspired various video-specific high-performing designs. For example, some works leverage extracted optical flow features [6, 14] or motion vectors [50, 51]. Others, like SlowFast [13], employ specific architectural designs to account for the fact that pixel values change less in the temporal dimension than in the spatial dimensions. In contrast to these approaches, our goal is to devise a generic framework for modeling multi-dimensional data. Consequently, Mamba-ND treats videos as simple 3D arrays of RGB pixels, making no additional assumptions about the temporal structure. This is crucial because in certain use cases such as weather forecasting, the third dimension of the data represents an additional spatial dimension instead of a temporal one.
@@
Climate Modeling.
Climate modeling presents a significant challenge in the 3D domain. We specifically focus on global weather forecasting. Typically, this task is evaluated using the ERA5 reanalysis dataset [18]. Early works [48, 49] employed CNN architectures. Pangu-Weather [3] was the first to introduce Transformers to this domain. It adapted a 3D-Swin Transformer to handle atmospheric measurements in 3D spaces. ClimaX [34] extended the Vision Transformer to weather data and proposed an architecture called Cli-ViT. Using this architecture, it established a strong foundational model, called ClimaX, capable of various tasks such as forecasting and downscaling through large-scale pretraining on simulated data.
Using state space models for climate modeling remains an unexplored area. We are the first to apply SSMs in the 3D weather forecasting task.
@@ 3.1 State Space Models
Early works such as S4 [16] and S4ND [33] assume linear time invariance (LTI). This constraint makes it possible to solve the above difference equation using a global convolution kernel. Selective state space models, i.e., Mamba [15], introduce time-varying parameters that do not follow the LTI assumption. In particular, $\Delta$, $B$, and $C$ become functions of the input signal $x(t)$. Consequently, from Equations 3, 4, 5 and 6, $\bar{A}$ and $\bar{B}$ also become dependent on time. This makes it hard to parallelize the computation. A hardware-aware optimization that makes use of associative scan is employed to mitigate this issue.
@@ 3.2 Mamba Layers
Mamba [15] proposed an implementation of selective state space model (sSSM) layer (Figure 3 column 1). It consists of a 1D convolution, an SSM kernel, and a residual connection. In each layer, the input sequence is first processed by the convolution operation and then by the SSM kernel. The result is added back to the input through the residual connection. While the Conv1D operation can be easily extended to multiple dimensions with ConvND operations, it is non-trivial to convert the SSM to multiple dimensions. S4ND [33] resolved this by formulating higher-dimensional SSM as a time-invariant system of partial differential equations (PDEs), which can be computed using global convolution. However, since Mamba is not a time-invariant system, this approach is not applicable. In this work, we discuss various alternative ways to extend SSM to higher dimensions by flattening the input data into 1D sequences.
@@ 4 Methodology
Figure 2: Architecture of Mamba-ND: In this figure we visualize Mamba-3D. Given 3D input data, we patchify it into $L$ patches. During this process, we maintain the original 3D structure of the input. This sequence is then passed through $K$ Mamba-ND blocks, each of which consists of a chain of 1-D Mamba layers that process the sequence in alternating orderings. In 3D space, we use the order H+H-W+W-T+T-. In 2D space, the sequence would be H+H-W+W-. Finally, the sequence is reshaped back to its original 3D structure and passed to task-specific heads for downstream tasks.
Mamba-ND explores several approaches that adapt Mamba to multidimensional data. The key element of these designs is to devise a combination of sequence orderings to flatten the multidimensional data into 1-D sequences. Intuitively, some level of bidirectional or multidirectional design is required to allow information exchange between two arbitrary data points in multidimensional space. This can be achieved at two levels: layer level and block level. As mentioned in Section 3.2, a Mamba layer consists of a 1D convolution, an SSM kernel, and a residual connection. One example of layer-level design is to pass the output of the convolution to two independent SSM kernels and sum up the results (Figure 3, col 2). Contrary to layer-level designs, block-level design keeps the internal design of a Mamba layer unchanged. Instead, it tries to achieve bidirectionality on the block level. For example, one block-level design is to alternate between each axis from one layer to another and apply a bidirectional design for each axis (Figure 6, row 2). One possible benefit of layer-level design is that it keeps the optimized fused kernel of a Mamba layer and the memory access pattern of the underlying array unchanged.
@@
Figure 5: Different ways of arranging Mamba layers. The first row visualizes alternating-directional design. The second row visualizes bidirectional design. The third row visualizes quad-directional design.
1
Figure 5: Different ways of arranging Mamba layers. The first row visualizes alternating-directional design. The second row visualizes bidirectional design. The third row visualizes quad-directional design.
Figure 6: Visualization of various Scan-Factorization policies. Col 1: No factorization, there is only 1 sequence. Col 2: Factorizing the 3D sequence into $D$ 2D sequences, where $D$ is the length of a single dimension. Col 3: Factorizing the 3D sequence into $D^{2}$ 1D sequences.
@@ 4.1 Scan Orderings
Consider the input N-dimensional data $X$ of shape $D_{1}\times D_{2}\times\ldots\times D_{N}$, where $D_{i}$ is the length of data along the $i$th dimension. Let $L=\prod_{i=1}^{N}D_{i}$ be the total sequence length. There are a total of $L!$ possible ways of flattening $X$ into a 1D sequence. However, we only consider a subset of these choices, which we call the scan orderings. Formally, a scan ordering is obtained by permuting the order of axes of the original data $X$, and flattening it into a 1D sequence either in the forward or reverse direction. Since there are $N!$ permutations of $N$ axes and 2 possible directions, there are a total of $2N!$ scan orderings.
We denote a particular scan ordering $s$ as $(k_{1},k_{2},\ldots,k_{N})\pm$, which represents the unique ordering obtained by first permuting the axis order from $1,2,\ldots,N$ to $k_{1},k_{2},\ldots,k_{N}$, and then flattening the sequence in row-major order. The symbols $+$ and -- indicate whether the order of the final 1D sequence is in forward or reversed direction.
@@
For simplicity, we use $H$ to represent $WH$ and $W$ to represent $HW$ for 2D data. Similarly, we use $H$ to represent $TWH$, $W$ to represent $THW$, and $T$ for $HWT$ for 3D data. In this notation, $(WH)-$ becomes $H-$ and $(HWT)+$ becomes $T+$. We also use $L$ to represent $THW$ or $HW$, as this is the naive way of flattening 3D and 2D data to a 1D sequence without changing the memory layout. Notably, the last dimension will be traversed continuously. We visualize 2D scan orderings in Figure 4.
@@ 4.2 Adapting the Mamba Layer
We explore three possible alterations to the standard Mamba layer design, which are illustrated in Figure 3.
Bi-SSM
layer passes the output of the convolution layer to two independent SSM kernels, one in the forward direction and another in the reversed direction. This design draws inspiration from bidirectional LSTM [7], in which two LSTM layers of opposite directions are stacked together in a similar fashion.
ND-SSM
layer extends Bi-SSM by incorporating additional SSMs to accommodate other possible orderings. In the 2D case, there are four orderings $W+$, $W-$, $H+$, $H-$.
Multi-head SSM
layer is a mimic of the multi-head design in the Transformer. It splits an input sequence of dimension $D$ into $H$ sequences of dimension $D/H$, where $H$ is the number of orderings. Each of the heads is then passed to separate SSM kernels in respective orderings. In the 2D case, the orderings are $W+$, $W-$, $H+$, $H-$.

### 12 Mar 2024 v2

@@ Abstract
In recent years, Transformers have become the de-facto architecture for sequence modeling on text and multi-dimensional data, such as images and video. However, the use of self-attention layers in a Transformer incurs prohibitive compute and memory complexity that scales quadratically w.r.t. the sequence length. A recent architecture, Mamba, based on state space models has been shown to achieve comparable performance for modeling text sequences, while scaling linearly with the sequence length. In this work, we present Mamba-ND, a generalized design extending the Mamba architecture to arbitrary multi-dimensional data. Our design alternatively unravels the input data across different dimensions following row-major orderings. We provide a systematic comparison of Mamba-ND with several other alternatives, based on prior multi-dimensional extensions such as Bi-directional LSTMs and S4ND. Empirically, we show that Mamba-ND demonstrates performance competitive with the state-of-the-art on various multi-dimensional benchmarks, including ImageNet-1K classification, HMDB-51 and UCF-101 action recognition, ERA5 weather forecasting and BTCV 3D segmentation. Code is available at https://github.com/jacklishufan/Mamba-ND
Keywords:
State Space Models Multi-Dimensional Modeling
@@ 1 Introduction
The design of flexible and scalable neural network architectures is fundamental to the success of deep learning across diverse domains. Convolutional neural networks [25] excel at handling continuous data such as images, audio, and video. Recently, they have been surpassed by Transformers [50], which process continuous data as a discrete sequence of patches [11]. Despite their superior performance on many tasks, Transformer-based models struggle to scale to larger patch sequence as they scale quadratically with respect to sequence length. Most recently, a special kind of State Space Model (SSM) termed as Mamba [14], has demonstrated stronger performance than Transformers while maintaining a linear complexity. However, the impressive performance of Mamba was shown on 1D text sequences. This leaves open the question whether Mamba can be effectively extended to multi-dimensional data such as images, video, or scientific datasets, which is the focus of this work.
@@
Unlike convolution or self-attention operations, which can be computed in parallel across the ND input data, Mamba requires a specific ordering of the data. Determining such an order is not an easy task. Building on past work in architecture design, we could consider many choices. One naive approach is to flatten the data in row-major order. Intuitively, this is non-optimal because in this setting information only flows in one direction, which is suboptimal for multi-dimensional data with no default ordering. Drawing inspiration from early works on Bi-directional LSTM, another alternative is to process the sequence in two directions at each layer and aggregate the results. We call this method Bi-SSM. While this design, in principle, allows information exchange between two arbitrary patches, two patches adjacent to each other spatially may have a huge distance between them on the computation graph. A natural extension of this method, which we call ND-SSM, is to process the input in $2D$ directions, where $D$ is the dimension of the data, and aggregate the results. Rather than picking orders in sequence, another possibility is to borrow inspiration from multi-head self-attention in transformers, wherein we can split the channels into multiple heads and let each head process an SSM in a different direction. This design is similar to ND-SSM but differs in that it has less computational burden per layer.
@@
At a high level, there is also the question of block-level design, which specifies how Mamba layers are organized. For example, using the vanilla Mamba layer as a black box, one can still apply the bi-directional or N-directional design by processing the same input through multiple layers with different directions.
@@
In this work, we conducted an extensive study on these possible design choices. Surprisingly, we find that simply alternating between three fixed row-major orderings is one of the best-performing strategies on both 2D and 3D data. Armed with these findings, we propose Mamba-ND, a surprisingly simple yet effective design to extend Mamba to multi-dimensional data. Mamba-ND does not introduce any complicated changes to 1D-SSM layers. By stacking 1D-SSM layers as black boxes and alternating the sequence order between each layer, Mamba-ND was able to surpass Transformer-based models on various tasks, including image classification, action recognition, and weather forecasting and 3D segmentation, with a lower parameter count. It also maintains a linear complexity with respect to input sequence length.
@@
  • We propose Mamba-ND , which extends SSM to higher dimension through simply alternating sequence odering across layers.
  • Compared with Transformers, our model is able to achieve stronger performance at much lower parameter count on a diverse set of tasks.
  • We conduct extensive ablation studies on various more complicated approaches that extend SSM to multi-dimensional inputs. We find that complicated designs do not necessarily translate into stronger performance.
@@ 2 Related Works
Modeling 2D Data :
Convolutional Neural Networks (CNNs) [25, 18, 21, 44, 22, 42] have historically been the state-of-the-art approach for image recognition tasks. The receptive field of CNNs only grows linearly with the model depth. This limitation restricts the ability of CNNs to effectively capture global context.
More recently, Vision Transformers (ViTs) [11] have emerged as a strong candidate for vision tasks. These models first divide an input image into a grid of discrete patches and then process them as a 1D sequence. While they provide a global receptive field through self-attention mechanisms at every layer, such advantages come with the cost of quadratic computational and memory complexity with respect to input length, making them challenging to scale.
To address this limitation, several works have proposed hybrid architectures that incorporate attention mechanisms into CNNs [52, 5, 26] or introduce hierarchical designs into transformers [29, 27, 7, 9, 10, 32, 45, 51, 55]. Other approaches, such as PixelRNN [48], attempt to apply recurrent models, which are inherently designed for linear complexity and global context, to images. There has also been considerable efforts to apply state space models to 2D images, such as 2D-SSM[2] and S4ND [37]. Our work falls within this final category, wherein we specifically explore ways to adapt selective state space models to images.
Modeling 3D Data :
Modeling videos has long presented a significant challenge. Earlier works [4, 46] primarily focused on extending 2D ConvNets into the 3D domain. More recently, 3D Transformers [1, 30, 36] have demonstrated superior performance. The unique characteristics of the temporal dimension in video data also inspired various video-specific high-performing designs. For example, some works leverage extracted optical flow features [6, 13] or motion vectors [53, 54]. Others, like SlowFast [12], employ specific architectural designs to account for the fact that pixel values change less in the temporal dimension than in the spatial dimensions. In contrast to these approaches, our goal is to devise a generic framework for modeling multi-dimensional data. Consequently, Mamba-ND treats videos as simple 3D arrays of RGB pixels, making no additional assumptions about the temporal structure. This is crucial because in certain use cases such as weather forecasting, the third dimension of the data represents an additional spatial dimension instead of a temporal one.
@@
In addition to videos, we also consider the 3D climate forecast and medical image segmentation task. In both field transformers [38, 3, 17, 16] have achieved successes. However they face similar scaling challenges as in typical vision problems.
@@ 3.1 State Space Models
Early works such as S4 [15] and S4ND [37] assume linear time invariance (LTI). This constraint makes it possible to solve the above difference equation using a global convolution kernel. Selective state space models, i.e., Mamba [14], introduce time-varying parameters that do not follow the LTI assumption. In particular, $\Delta$, $B$, and $C$ become functions of the input signal $x(t)$. Consequently, from Eqs. 3, 4, 5 and 6, $\bar{A}$ and $\bar{B}$ also become dependent on time. This makes it hard to parallelize the computation. A hardware-aware optimization that makes use of associative scan is employed to mitigate this issue.
@@ 3.2 Mamba Layers
Mamba [14] proposed an implementation of selective state space model (sSSM) layer (Fig. 3 column 1). It consists of a 1D convolution, an SSM kernel, and a residual connection. In each layer, the input sequence is first processed by the convolution operation and then by the SSM kernel. The result is added back to the input through the residual connection. While the Conv1D operation can be easily extended to multiple dimensions with ConvND operations, it is non-trivial to convert the SSM to multiple dimensions. Particularly, the convolution trick is not applicable since Mamba is not a time-invariant system. In this work, we discuss various alternative ways to extend SSM to higher dimensions by flattening the input data into 1D sequences.
@@ 4 Methodology
Figure 2: Mamba-ND Architecture. We visualize Mamba-3D as an example. Given 3D input, we patchify it into $L$ patches. During this process, we maintain the original 3D structure of the input. This sequence is then passed through $K$ Mamba-ND blocks, each of which consists of a chain of 1D Mamba layers that process the sequence in alternating orderings. In 3D space, we use the order H+H-W+W-T+T-. In 2D space, the sequence would be H+H-W+W-. Finally, the sequence is reshaped back to its original 3D structure and passed to task-specific heads for downstream processing.
We explores several approaches to adapt Mamba to multidimensional data. The key element of these designs is to devise a combination of sequence orderings to flatten the multidimensional data into 1D sequences. Intuitively, some level of bidirectional or multidirectional design is required to allow information exchange between two arbitrary data points in multidimensional space. This can be achieved at two levels: layer level and block level. As mentioned in Sec. 3.2, a Mamba layer consists of a 1D convolution, an SSM kernel, and a residual connection. One example of layer-level design is to pass the output of the convolution to two independent SSM kernels and sum up the results (Fig. 3, col 2). Contrary to layer-level designs, block-level designs keep the internal design of a Mamba layer unchanged. Instead, it tries to achieve bidirectionality at the block level. For example, one such design is to alternate between each axis from one layer to another and apply a bidirectional design for each axis (Fig. 4(a), row 2). One possible benefit of layer-level design is that it keeps the optimized fused kernel of a Mamba layer and the memory access pattern of the underlying array unchanged.
@@
(a) Different ways of arranging Mamba layers. The first row visualizes alternating-directional design. The second row visualizes bidirectional design. The third row visualizes quad-directional design.
(b) Visualization of various Scan-Factorization policies. Col 1: No factorization, there is only 1 sequence. Col 2: Factorizing the 3D sequence into $D$ 2D sequences, where $D$ is the length of a single dimension. Col 3: Factorizing the 3D sequence into $D^{2}$ 1D sequences.
Figure 5: Visualization of block level design and factorization policies.
@@ 4.1 Scan Orderings
Consider the input N-dimensional data $X$ of shape $D_{1}\times D_{2}\times\ldots\times D_{N}$, where $D_{i}$ is the length of data along the $i$th dimension. Let $L=\prod_{i=1}^{N}D_{i}$ be the total sequence length. There are a total of $L!$ possible ways of flattening $X$ into a 1D sequence. However, we only consider a subset of these choices, which we call the scan orderings. Formally, a scan ordering is obtained by permuting the order of axes of the original data $X$, and flattening it into a 1D sequence either in the forward or reverse direction. Since there are $N!$ permutations of $N$ axes and 2 possible directions, there are a total of $2N!$ scan orderings. We denote a particular scan ordering $s$ as $(k_{1},k_{2},\ldots,k_{N})\pm$, which represents the unique ordering obtained by first permuting the axis order from $1,2,\ldots,N$ to $k_{1},k_{2},\ldots,k_{N}$, and then flattening the sequence in row-major order. The symbols $+$ and -- indicate whether the order of the final 1D sequence is in forward or reversed direction.
@@
For simplicity, we use $H$ to represent $WH$ and $W$ to represent $HW$ for 2D data. Similarly, we use $H$ to represent $TWH$, $W$ to represent $THW$, and $T$ for $HWT$ for 3D data. In this notation, $(WH)-$ becomes $H-$ and $(HWT)+$ becomes $T+$. We also use $L$ to represent $THW$ or $HW$, as this is the naive way of flattening 3D and 2D data to a 1D sequence without changing the memory layout. Notably, the last dimension will be traversed continuously. We visualize 2D scan orderings in Fig. 4.
@@ 4.2 Adapting the Mamba Layer
We explore three alterations to the standard Mamba layer design, which are illustrated in Fig. 3.
Bi-SSM layer passes the output of the convolution layer to two independent SSM kernels, one in the forward direction and another in the reversed direction.
ND-SSM layer extends Bi-SSM by incorporating additional SSMs to accommodate other possible orderings. In the 2D case, there are four orderings $W+$, $W-$, $H+$, $H-$.
Multi-head SSM layer is a mimic of the multi-head attention. It splits an input sequence of dimension $D$ into $H$ sequences of dimension $D/H$, where $H$ is the number of orderings. Each of the heads is then passed to separate SSM kernels in respective orderings. In the 2D case, the orderings are $W+$, $W-$, $H+$, $H-$.